Find the equations of the lines through the following pairs of points, in slope intercept form.
(1) Find the equation of the line through (2,1), (−6,8).

Answer by texttutoring(324)

First we need to find the slope of the line connecting those 2 points.
The formula for slope is:
m=(y2-y1)/(x2-x1)
In your question, x1=2,y1=1 and x2=-6,y2=8
m= (8-1)/(-6-2)
m= 7/-8
Now we can use the equation of a line (y=mx+b) to find the y-intercept (b).
We can choose either point (2,1) or (-6,8) to plug in for x and y. I will choose (2,1) because it is easier.
y=mx+b
1=(-7/8)*2+b
1=-14/8 + b
8/8+14/8=b
b=22/8
In slope-intercept form, the equation of the line is:
y=(-7/8)x + 22/8